OpenStudy (anonymous):

find all polar coordinates of point p where p = (9, -pi/5)

OpenStudy (anonymous):

i thought we did this one?

OpenStudy (anonymous):

i know but it wasnt right

OpenStudy (anonymous):

they are asking for two

OpenStudy (anonymous):

you can add any multiple of \(2\pi\) to the angle

OpenStudy (anonymous):

for example

OpenStudy (anonymous):

it says "all"

OpenStudy (anonymous):

(9, (9, -pi/5) + 2nπ) or (9, -(9, -pi/5) + (2n + 1)π)

OpenStudy (anonymous):

i guess you could also say \[(-9,\frac{4\pi}{5})\]

OpenStudy (anonymous):

This is one of the answers (9, (9, -pi/5) + 2nπ) or (9, -(9, -pi/5) + (2n + 1)π)

OpenStudy (anonymous):

i am afraid i don't understand what that is supposed to mean

OpenStudy (anonymous):

These are the list of answers to the question (9, -pi/5 + 2nπ) or (-9, pi/5 + 2nπ) (9, -pi/5 alt tag: negative pi divided by 5 + (2n + 1)π) or (-9, pi/5 + 2nπ) (9, -pi/5 + 2nπ) or (-9, pi/5 + (2n + 1)π) (9, -pi/5 + 2nπ) or (9, pi/5 + (2n + 1)π)

OpenStudy (anonymous):

@phi can u help me

OpenStudy (anonymous):

ahh this makes more sense this one (9, -pi/5 + 2nπ) works for sure

OpenStudy (anonymous):

their asking for the answer to be in a pair

OpenStudy (anonymous):

I think its this one (9, -pi/5 + 2nπ) or (9, pi/5 + (2n + 1)π)

OpenStudy (anonymous):

yes that one is it kind of a weird way to write it

OpenStudy (anonymous):

i know but thx

OpenStudy (anonymous):

oh wait, let me check

OpenStudy (phi):

starting with (9, -pi/5) you have (9, -pi/5 + 2 n pi) (-9, -pi/5 + pi + 2 n pi) or (-9, -pi/5 - pi + 2 n pi) (if you change 9 to -9 and add (or subtract) pi to the angle, you are back where you started) simplifying (-9, -pi/5 + pi + 2 n pi) you get (-9, -pi/5 + (2n+1) pi ) simplifying (-9, -pi/5 - pi + 2 n pi) you get (-9, -pi/5 + (2n-1) pi ) I think you want (9, -pi/5 + 2nπ) or (-9, -pi/5 + (2n + 1)π) but I can't tell if that is one of your choices.